HUBBLE SPACE TELESCOPE DAILY REPORT #2474 PERIOD COVERED: 0000Z (UTC) 10/07/99 - 0000Z (UTC) 10/08/99 Daily Status Report as of 281/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 1.3 Completed STIS/CCD 8313 (Imaging the Host Galaxies of High Redshift Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of 974 and 9748 and their host galaxies. The propose is to obtain snapshot, unfiltered STIS images of distant galaxies of known redshift which have hosted supernovae {SNe} of Type IA found via the Supernova Cosmology Project. The observations were completed as scheduled, and no anomalies were reported. 1.4 Completed Two Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.5 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.6 Completed WF/PC-2 6752 (Global Mapping of the Opposition Surge on the Galilean Satellites) The WF/PC-2 was used to make observations of Jupiter's moons Ganyemede, Io, Europa and Callisto. This is to construct ~200 km resolution global maps of the opposition brightness surge for each of the four Galilean satellites to support and complement Galileo mission science investigations. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/MA1 8428 (STIS FUV-MAMA Cycle 8 Flats) The Space Telescope Imaging Spectrograph (MA1) was used to make a series of far-ultraviolet calibration flats with the Kr lamp for the construction of on-orbit D-flats for select modes. The calibration images were completed as planned, and no problems were reported. 1.8 Completed WF/PC-2 8243 (SINS: The Supernova Intensive Study - Cycle 8) The WF/PC-2 was used to make observations of supernova LMC-SN1987. Supernovae reveal the last stages of stellar evolution, create the chemical history of the Universe, energize the interstellar gas, and occupy the center of attention for the extragalactic distance scale and cosmology. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.9 Completed STIS/CCD/MA2 8257 (VV Cephei: The Egress from Chromospheric Eclipse) The Space Telescope Imaging Spectrograph was used to make choromospheric observations of the egress from eclipse of star HD60414. This is to make observations of the long-period {20.3 yrs} eclipsing binary VV Cephei {M2 Iab + B0} as it continues to emerge from chromospheric eclipse. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 8390 (Expansion Parallax Distances to Planetary Nebulae) The WF/PC-2 was used to make observations of the planetary nebula IC2448 in order to study its parallax distance. The largest impediment to our understanding of planetary nebulae is the lack of reliable distances. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Four Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed STIS/CCD 8549 (POMS Test Proposal: STIS non-scripted parallel proposal continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations as part of a non-scripted POMS Test Proposal. The observations were completed as scheduled, and no problems were reported. 1.14 Completed STIS/CCD 8229 (Spatially Resolved Spectroscopy of Titan's Atmosphere and Surface) The Space Telescope Imaging Spectrograph (CCD) was used to observe the moon Titan. The observations will measure the surface reflectivity of Titan at 0.9 microns, and map the distribution of haze in the northern and southern hemispheres at altitudes where the haze is formed. The proposal completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 8 Scheduled Re-acquisitions: 7 Successful: 7 2.2 FHST Updates: Scheduled: 22 Successful: 22 2.3 Operations Notes: The SSR EDAC error counter was cleared three times per ROP SR-1. A TTR was written since an NCC cold start resulted in no GCMR capability from 280/1710Z until 280/1721Z. The NICMOS FPA temperature sensor for detector 3 was enables at 280/2133Z. The CCS FEP did not ingest data for three times slots: from 281/0049Z-0127Z, from 281/0144Z-0212Z, and from 281/0231Z-02238Z. At the same times PRS showed no problems, and was locked on data. The response team was notified. The problem did not re-occur. The cause is currently unknown. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH